OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (1, 9) and (−1, 11)? y + 9 = −(x − 1) y − 9 = 1(x + 1) y + 9 = 1(x + 1) y − 9 = −(x − 1)

OpenStudy (anonymous):

ok nevermind those.im half asleep

OpenStudy (anonymous):

point slope is y-y1=m(x-x1)

OpenStudy (anonymous):

am i right @skullpatrol

OpenStudy (skullpatrol):

$$\huge y-y_1=m(x-x_1)$$

OpenStudy (anonymous):

first you need to find the slope

OpenStudy (anonymous):

so what is the slope of the line @BIZZLE05

OpenStudy (skullpatrol):

The formula is right @marinesrock but I prefer the student to ask MORE questions :)

OpenStudy (anonymous):

This is the formula for that

OpenStudy (skullpatrol):

Without the student putting in effort to ask questions, no amount of teaching will help.

OpenStudy (anonymous):

ur right @skullpatrol i will stand by and if he makes a mistake i will point it out

OpenStudy (anonymous):

i don't know it, that's why i asked the question

OpenStudy (anonymous):

ok make a graph and graph the points.

OpenStudy (skullpatrol):

That @BIZZLE05 is not a question, it's request for the answer.

OpenStudy (skullpatrol):

What makes math so hard is that to understand it you have to do it yourself.